# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

### v1

@@ 5.2. Full Confluence Checking using Cut Points
Figure 5. Evaluation results on a benchmark of 20 programs. From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.). All test cases are written originally in C and then compiled to LLVM via Clang (clang, ). The first 10 programs are benchmark programs from the RipTide work. The nn_* tests implement neural network inference. The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ). In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug. In all other cases, the bisimulation and confluence checks succeed.
Name LOC #OP #PC Bisim. (sec) Conf. (sec)
bfs 48 48 3634 0.45 7.79
dconv 24 65 8134 0.96 11.87
dfs 45 49 3762 0.46 8.51
dmm 16 56 4220 0.71 6.31
dmv 10 31 1234 0.22 1.69
fft 29 70 2690 0.45 6.04
sconv 22 56 4272 0.71 10.89
smm 22 49 3967 0.53 12.08
smv 12 25 1103 0.19 2.36
sort 23 57 6554 $\times$ 0.64 3.62
nn_conv 43 61 12154 0.78 15.65
nn_fc 24 36 2632 0.31 3.43
nn_norm 15 19 316 0.07 0.17
nn_pool 33 52 8165 0.74 7.11
nn_relu 16 19 844 0.11 0.44
nn_vadd 9 15 244 0.06 0.15
Dither 21 31 2553 0.32 2.36
SpMSpMd 42 60 7693 0.93 28.45
SpMSpVd 48 40 3182 0.4 10.22
SpSlice 33 31 2616 $\times$ 0.34 7.4
@@
Figure 5. Evaluation results on a benchmark of 20 programs. From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.). All test cases are written originally in C and then compiled to LLVM via Clang (clang, ). The first 10 programs are benchmark programs from the RipTide work. The nn_* tests implement neural network inference. The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ). In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug. In all other cases, the bisimulation and confluence checks succeed.
@@ 6. Implementation and Evaluation
We have implemented our translation validation technique in the tool WaveCert  [^2] Available on GitHub at https://github.com/secure-foundations/riptide-verification.. We first implemented symbolic executors for (a subset of) LLVM and dataflow programs in Python (z3, ). Based on these, we implemented the bisimulation check in Section 4 and the confluence check in Section 5.
WaveCert uses the Z3 SMT solver (z3, ) to discharge the verification conditions it generates. In the bisimulation check, WaveCert encodes the validity of the correspondence conditions at each cut point into Z3 queries. In the confluence check, WaveCert produces a list of permission constraints and checks if these constraints are satisfiable with respect to the finite permission algebra (Definition 2) by encoding each permission variable as a set of Boolean variables and the list of permission constraints as a SAT query.
In order to evaluate WaveCert’s capability to certify compilation to real dataflow programs, we have applied WaveCert to a benchmark of 20 programs consisting of the benchmark programs in RipTide (riptide, ) and Pipestitch (pipestitch, ), as well as some programs implementing neural network inference. Figure 5 shows some statistics about the benchmark programs and how WaveCert performs. All tests are performed on a laptop with an M1 Max processor and 64 GiB of RAM.
In all of the benchmark programs, the bisimulation check either succeeds or finds an actual bug (such as in sort and SpSlice, which we discuss further in Section 6.1). This shows that our heuristics for inferring dataflow cut points works well on the canonical schedule. Our bisimulation check generates verification conditions in quantifier-free theories of arrays and integers without non-linear arithmetic, so all bisimulation checks finish quickly within a second.
The confluence check succeeds in all examples. For sort, we manually fixed the incorrect dataflow program produced by the compiler, after which the confluence check passes.
@@
Figure 6. Examples of two compiler bugs.
@@ 6.1. Compiler Bugs
During the testing of WaveCert, we found 7 bugs in the RipTide compiler. Three of these bugs were discovered on the test cases sort, SpSplice, and their variations; and the rest four bugs were found on smaller test cases that we manually constructed. A developer of the RipTide compiler has confirmed these issues.
In the following, we discuss two examples of these bugs to show that the bisimulation check and the confluence check are both effective in catching compilation bugs.
@@
WaveCert found a compilation bug involving integer signed vs. unsigned extension in the sort test case. A simplified LLVM program used to reproduce this issue is shown in the top left of Figure 6, and the output dataflow program is shown in the top right of Figure 6. The issue itself is straightforward: in the LLVM code, a 1-bit xor is done with a 1-bit constant 1, and then the result is zero-extended to a 32-bit integer. On the dataflow side, the dataflow compiler tries to convert all operations into 32-bits, the word width supported on the RipTide architecture. However, during the process, the compiler assumes that all constants should be sign-extended. As a result, it extends the 1-bit 1 to a 32-bit -1, which is incorrect in this case. Our bisimulation check catches this bug when verifying the correspondence condition at the final cut point, where the memory configurations do not match.
@@
Our confluence check catches a compilation issue when compiling from the C program shown in the bottom left of Figure 6 to the dataflow program in the bottom right of Figure 6. In the C program, we have a nested loop. In each iteration of the outer loop, we first set A[i] = i, and then in the inner loop that follows, we copy A[i] into A[i + 1], …, A[n - 1]. In the output dataflow program, the compiler allows the load and store operators in the inner loop to pipeline using the invariant operator (marked I) to repeatedly send the signal to enable load. However, this causes a data race, because in the inner loop, a load in the second iteration accesses the same memory location as the store in the first iteration.
Our confluence check fails as the list of permission constraints we collected are unsatisfiable. WaveCert also outputs an unsat core of 57 constraints, a subset of the 420 constraints generated, that causes the unsatisfiability. In particular, the unsat core includes two ordering constraints: read  A is contained in the disjoint union of input tokens of the inner loop load and write  A is contained in the disjoint union of input tokens of the inner loop store. This indicates that the main conflict stems from the inner loop load and store.
@@ 7. Limitations and Future Work
Our translation validation technique still has some limitations, and we plan to address them in future work.
@@
The first two optimizations are not supported by our bisimulation check because we require a one-to-one correspondence between LLVM instructions and dataflow operators for the canonical schedule. In future work, we can relax this requirement and allow multiple LLVM instructions to map to one dataflow operator, and we then only fire the operator if all of the corresponding LLVM instructions have fired.
@@
Liveness properties
Our semantics of the dataflow program uses unbounded FIFO queues to model dataflow channels. In this model, we do not need to consider liveness issues such as deadlocks, because an operator can always fire once the input values are ready. However, this is not true for the actual hardware, which has a limit on the queue size.
For future work, we plan to explore model checking (model-checking, ) techniques to check for deadlocks.
@@ 8. Related Work
As a compiler verification technique, the translation validation approach taken in our work is a less labor-intensive alternative to full compiler verification, such as CompCert (compcert, ). It was first proposed by Pnueli et al. (pnueli-tv, ), and there are numerous following works with novel translation validation or program equivalence checking techniques for various languages and compilers (necula-tv, ; kundu-tv, ; data-driven-eq, ; tate-eq-saturation, ). For LLVM-related works on translation validation, we have Alive (alive, ; alive2, ), which performs bounded translation validation for internal optimizations in LLVM. LLVM-MD (llvm-md, ; denotational-tv, ) is another translation validation tool for LLVM optimizations. LLVM-MD translates the LLVM code into an intermediate representation called synchronous value graphs (SVG) (denotational-tv, ), which is a similar representation to a dataflow program. The difference is that the semantics of SVG is synchronous, and they also trust the compilation from LLVM to SVG, whereas in our case, such translation is exactly what we are trying to validate. Our formulation of cut-bisimulation in our bisimulation check is an adaptation from the work on Keq (keq, ), which aims to be a more general program equivalence checker parametric in the operational semantics of input/output languages.
Our use of linear permission tokens is a dynamic variant of the fractional permissions proposed by Boyland (fractional-perm, ) and used in some variations of concurrent separation logic (concurrent-separation-logic, ). In contrast to separation logic frameworks such as Iris (iris, ), which are embedded in interactive proof assistants, our permission tokens are synthesized and checked fully automatically.
There are also verification tools for classical dataflow languages such as Lustre (lustre, ; lustre-verified-compiler, ; lustre-smt-verification, ) and Esterel (esterel-verification, ). These languages have a synchronous semantics and do not interact with globally accessible memory, whereas dataflow programs emitted by CGRA compilers are highly asynchronous and make pervasive use of the global heap.
@@
In the context of term rewriting systems (term-rewriting, ), the standard approach for confluence checking is the use of critical pairs (knuth-critical-pairs, ; duran-confluence, ), which proves ground confluence in the case when the rewriting system is terminating. This algorithm is implemented in various systems, such as the Church-Rosser checker in Maude (maude, ). However, this technique is difficult to apply in our case, since the semantics of a dataflow program is neither terminating nor, in general, ground confluent.
@@ 9. Conclusion
In this work, we develop a translation validation technique for asynchronous dataflow programs with an application to the RipTide CGRA architecture. Our technique first checks the bisimulation of the input imperative program and the output dataflow program on a canonical schedule, and then we generalize the results to all possible schedules via linear permission tokens. Our work is an initial step towards integrating high assurance into new architectures from the start, rather than retrofitting them decades later.

### v2

@@ 5.3. Liveness in the Bounded-Channel Model
Figure 5. Evaluation results on 21 test cases sorted by the number of dataflow operators. From left to right, the columns are the name of the test case (Name), lines of code in the original C program (LOC), number of dataflow operators (#OP), number of permission constraints (#PC), time spent for simulation check in seconds (Sim.), and time spent for confluence check in seconds (Conf.). All test cases are originally written in C and then compiled to LLVM via Clang (LLVM, 2024a). Compiler bugs cause the simulation checks to fail in SpSlice, sort, and sha256. In all other cases, the simulation and confluence checks succeed.
[…]
@@ 6. Implementation and Evaluation
We have implemented our translation validation technique in the tool WaveCert targeting the RipTide compiler from LLVM programs to dataflow programs. It is publicly available in a GitHub repository (Lin et al., 2024). We implemented symbolic executors for (a subset of) LLVM and dataflow programs in Python. Based on these, we implemented the cut-simulation check described in Section 4 and the confluence check in Section 5. We also instrumented the RipTide compiler to output hints for constructing the canonical schedule (Section 4).
WaveCert uses the Z3 SMT solver (De Moura and Bjørner, 2008) to discharge the verification conditions it generates. In the simulation check, WaveCert encodes feasibility of path conditions and validity of the correspondence conditions at each cut point into Z3 queries. In the confluence check, WaveCert produces a list of permission constraints and checks if these constraints are satisfiable with respect to the finite permission algebra (Definition 3) by encoding each permission variable as a set of Boolean variables and the list of permission constraints as an SMT query.
In order to evaluate WaveCert’s capability to certify compilation to real dataflow programs, we have applied WaveCert to a benchmark of 21 programs consisting of the benchmark programs in RipTide (Gobieski et al., 2023) and Pipestitch (Serafin et al., 2023), as well as some programs implementing neural network inference and an implementation of SHA-256 (Conte, 2024). Figure 5 shows some statistics about the benchmark programs and how WaveCert performs. All tests are performed on a laptop with an Apple M1 processor and 64 GiB of RAM. The time spent in the RipTide compiler (including hint generation) is less than 0.1 seconds for all of the benchmark programs, so it is omitted from the table.
We now evaluate WaveCert in three aspects: generality of cut-point heuristics, performance, and size of the trusted computing base.
Generality of the heuristics
Both simulation and confluence checks pass for most of the benchmark programs. This shows that our heuristics for inferring dataflow cut points work well on the canonical schedule, and our confluence check is general enough for common compilation patterns. In test cases sort, SpSlice, and sha256, the simulation check fails and correctly reveals two compilation bugs (see Section 6.1 for more detail). After fixing these bugs, the simulation and confluence checks succeed for these two programs.
Performance
WaveCert spends most of its time on the confluence check, mainly consisting of checking satisfiability for permission constraints. The simulation check takes a relatively short amount of time, because the input LLVM program and the output dataflow program match well on the canonical schedule and thus the verification conditions are simple to solve. Overall, the verification time is within 15 seconds for most benchmark programs, and we believe it is feasible to enable WaveCert in real production scenarios.
Trusted Computing Base
The original RipTide compiler has ~21,000 lines of C++ code, excluding comments. The instrumented hint generation in the RipTide compiler is implemented in less than 50 lines of C++. Compared to the RipTide compiler, the trusted computing base of WaveCert has 3,121 lines of Python with three main components: ~800 lines of dataflow semantics; ~1,000 lines of LLVM semantics; and ~750 lines for simulation and confluence checks.
@@
Figure 7. Deadlock example.
[…]
Deadlock example.
@@ 6.1. Compiler Bugs
While testing WaveCert, we found 8 bugs in the RipTide compiler, confirmed by RipTide authors. Four of these bugs were discovered on the test cases sort, SpSplice, sha256, and their variations; and the remaining four bugs were found on test cases not included in the evaluation. We now discuss two representative bugs:
@@
WaveCert found a compilation bug involving integer signed vs. unsigned extension in the sort test case. A simplified LLVM program used to reproduce this issue is shown in the top left of Figure 7, and the output dataflow program is shown in the top right of Figure 7. The LLVM code performs a 1-bit xor with a 1-bit constant 1, and it zero-extends the result to a 32-bit integer. On the dataflow side, the dataflow compiler converted all operations into 32-bits (word length of RipTide). However, during the process, the compiler assumes that all constants should be sign-extended. As a result, it extends the 1-bit 1 to a 32-bit -1, which is incorrect in this case. This difference causes our simulation check to fail, as the final memory states fail to match.
@@
Our confluence check catches a compilation bug when compiling from the C program shown in the bottom left of Figure 7 to the dataflow program in the bottom right of Figure 7. The C program has a nested loop. In each iteration of the outer loop, it first sets A[i] = i, and then in the inner loop that follows, it copies A[i] into A[i + 1], …, A[n - 1]. In the output dataflow program, the compiler allows the load and store operators in the inner loop to pipeline using the invariant operator (marked I) to repeatedly send the signal to enable load. However, this causes a data race, because in the inner loop, a load in the second iteration accesses the same memory location as the store in the first iteration.
The confluence check fails as the generated permission constraints are unsatisfiable. WaveCert also outputs an unsat core of 57 constraints, a subset of the 420 constraints generated, that causes the unsatisfiability. In particular, the unsat core includes two constraints: read  A is in the input tokens of the inner loop load, and write  A is in the input tokens of the inner loop store. This indicates that the main conflict stems from these two operators.
@@ 7. Limitations and Future Work
Our translation validation technique still has some limitations to address in future work.
@@
The first two optimizations are not supported by our simulation check because we require a one-to-one correspondence between LLVM instructions and dataflow operators for the canonical schedule. Since these optimizations involve reasoning about the equivalence between dataflow programs, we believe that a modular approach is to use WaveCert to verify the input LLVM against the unoptimized dataflow program, and use another verification pass in future work to soundly optimize the dataflow program.
@@
Scalability of the confluence check
WaveCert currently spends most of its time on the confluence check (Section 6). We believe this is primarily due to a naive encoding of permission constraints in SMT. For better performance in future work, one could design a custom solver for permission constraints to better utilize the structure of a permission algebra, or use a preprocessing step to simplify permission constraints.
@@ 8. Related Work
As a compiler verification technique, the translation validation approach taken in our work is a less labor-intensive alternative to full compiler verification, such as CompCert (Leroy, 2009). Translation validation was first proposed by Pnueli et al. (1998), and there are numerous following works with novel translation validation or program equivalence checking techniques for various languages and compilers (Necula, 2000; Kundu et al., 2009; Sharma et al., 2013; Tate et al., 2009). For LLVM-related works on translation validation, we have Alive (Lopes et al., 2015, 2021), which performs bounded translation validation for internal optimizations in LLVM. LLVM-MD (Tristan et al., 2011; Govereau, 2012) is another translation validation tool for LLVM optimizations. LLVM-MD translates the LLVM code into an intermediate representation called synchronous value graphs (SVG) (Govereau, 2012), which is a similar representation to a dataflow program. The difference is that the semantics of SVG is synchronous (in the sense that values are not buffered in the channels), and they also trust the compilation from LLVM to SVG, whereas in our case, such translation is exactly what we are trying to validate. Our formulation of cut-simulation in our simulation check is an adaptation from the work on Keq (Kasampalis et al., 2021), which aims to be a more general program equivalence checker parametric in the operational semantics of input/output languages.
Our use of linear permission tokens is a dynamic variant of the fractional permissions proposed by Boyland (Boyland, 2013) and used in various separation logics (Brookes, 2006) to reason about read-only sharing of references. Tools based on separation logics such as Iris (Jung et al., 2018) and Viper (Müller et al., 2016) usually require manual annotations for permission passing in the pre-/post-conditions, whereas our permission tokens are synthesized fully automatically. Furthermore, separation logic is not a suitable formalism to directly reason about dataflow programs due to the lack of structured control-flow in a dataflow program and the asynchrony of dataflow operators.
Various models of dataflow programming (Johnston et al., 2004) have been studied, such as computation graphs (Karp and Miller, 1966) and Kahn process networks (Gilles, 1974). An important aspect of these works is determinacy; i.e., these models produce a deterministic result regardless of the schedule of execution. There are also verification tools for classical dataflow languages such as Lustre (Halbwachs et al., 1991; Bourke et al., 2017; Hagen and Tinelli, 2008) and Esterel (Bouali, 1998). However, dataflow programs used by CGRA compilers often make use of shared global memory, which requires stronger techniques for guaranteeing determinacy.
Process calculi such as CSP (Roscoe, 1997) and $\pi$-calculus (Milner, 1999) present an alternative formalism for message-passing concurrency. To manage the inherent complexity and nondeterminism in process calculi, session types can enforce confluence and deadlock-freedom by construction (Sangiorgi and Walker, 2001). However, similar to separation logic, session types require manual typing annotations that would be difficult to synthesize automatically, and we likely still need some form of fractional permissions to handle shared memory.
Our strategy for analyzing nondeterministic dataflow programs is reminiscent of verification strategies for distributed systems (Bakst et al., 2017; Kragl et al., 2020; von Gleissenthall et al., 2019) which analyze message-passing protocols (e.g., Paxos (Lamport, 2001)) via reductions to sequential programs. However, the dataflow domain requires significantly different techniques: we verify equivalence properties with arbitrary specification programs, and use linear permissions to enable safe memory accesses in the presence of asynchrony.
@@
In the context of term rewriting systems (Baader and Nipkow, 1998), the standard approach for confluence checking is the use of critical pairs (Knuth and Bendix, 1970; Durán et al., 2020), which proves ground confluence in the case when the rewriting system is terminating. This algorithm is implemented in various systems, such as the Church-Rosser checker in Maude (Clavel et al., 2002). However, this technique is difficult to apply in our case, since the semantics of a dataflow program is neither terminating nor, in general, ground confluent.
@@ 9. Conclusion
In this work, we develop a technique for translation validation between imperative source programs and asynchronous dataflow target programs, with an implementation for the RipTide CGRA architecture. Our technique simplifies analysis on dataflow by first verifying that the target program simulates the source program along a canonical schedule, and then using linear memory permissions to show that this canonical schedule is general. Our verification procedure ensures both functional correctness and liveness for the target dataflow program.
In future work, we aim to support optimizations on dataflow programs (e.g., special-purpose operators for pipelining loops), and extending our verification methodology to concrete instantiations of CGRAs in hardware.